The vertex of parabola is at the point (-1, -3). how do i get the equation of the parabola?
Answer by stanbon(48581)

The vertex of parabola is at the point (-1, -3). how do i get the equation of the parabola?
------------------
Form y = ax^2 + bx + c
---------
The vertex occurs where x = -b/2a
So, -b/2a = -1 ; b = 2a
----------------------------------
Therefore, let a = 1,then b = 2
y = x^2 + 2x + c
-------
But y = -3 when x = -1,so:
-3 = 1 -2 + c
c = -2
----------------
EQUATION:
y = x^2 + 2x -2 is a parabola with vertex = (-1,-3).
==============================
Cheers,
Stan H.
====